Shift report 20000702 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sun Jul 02 2000 - 11:03:09 EDT

  • Next message: http account: "Shift report 20000702 08:00-16:00"

    ------------------------------------------------------------
    Date: Sun Jul  2 11:03:08 2000, Shift: 08:00-16:00
    Supervisor: fv, with: 
    Log book#2 pages 103 - 105
    
    Shift Summary:
    ------------------------------------------------------------
    NOTE: report is from July 1  08:00-17:30
    
    Beam was availbale in two periods after an early quenchlink.
    The quenchlink gave an opportunity for an access.
    
    MTP1 driftvoltage hv cable was moved to channel 7.2 after it was verified that the problem was external to TPC.Since then it has worked well.
    
    The nimbin 3 in rack 3 went into the mode where 24V drops to about 3.5. This time it would not go away with the Philips 794 GG in place. The GG (sync and pulser trigger)
    was moved to another crate. The Nim bin should be checked ouy.
    
    The fan problem reported in the shift check during two days
    was simple to fix. It was not plugged into the power strip.
    
    Part of shift was spend in introducing counting house procedures to new people.
    
    After inspection of the previous nights run it was realized that trigger 4 (ZDC) had been left (by me) from about 2 days ago during timing setup in single not coincidence mode.
    
    
    Once beam was back the collision rate was low (one per 5-10 sec). Despite that the time was used to checkout various trigger timing. It was confirmed that true ZDC triggers were
    about 50-60 nsec later that beam collision trigger. From the timing and cable layout it should be possible to have
    the ZDC in at LVL0.
     Therefore this time diff was reduced in the following way.
    a) tightened up of ZDC trigger cable path
    b) adding about 28-32 nsec to other triggers
    
    To adjust gate timing the following steps was done so far
    - removed 16 nsec for FB gate distribution using a short cable
    - removed 32 nsec from TDCs to BB, H1 and H2
    - still to be checked TOFW TDCs as well as ADC gates with collision trigger, not with the background as has mainly been done.
    
    The runs from Saturday afternoon on is with this setup.
    The data should be checked carefully.
    
    
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sun Jul 02 2000 - 11:07:19 EDT